Hi i am a high school student! my question is
we have a gas and we dont care about temperature change. the pressure is given from the type P(V)=6/V and the starting V is Vo. what will happen to the pressure if V is increased by 8%.
I know from the theory that it should decrease by 8% due to boyle's law BUT my maths fail me and i cannot prove it
1
Expert's answer
2020-08-31T12:46:29-0400

Hi!

Just write the law:


"P=\frac{6}{V}."

You said the pressure is increased by 8% compared to the initial pressure "P_0", which means


"P=\frac{(100\%+8\%)}{100\%}P_0=1.08P_0."

What will happen to the volume? From the very first equation express volume by multiplying both sides of the equation by V and dividing both sides by P:


"V=\frac{6}{P}."

Substitute our new pressure for P:


"V=\frac{6}{1.08P_0}=\frac{5.56}{P_0}."

Now compare how is this different from initial volume, which is just 6/P?


"\frac{V}{V_0}=\frac{5.56/P_0}{6/P_0}=0.926."

In other words (multiply both sides if this equation by V0):


"V=0.926V_0."

So, by how many percent is V smaller than V0? Let's calculate this.


Find the difference between the initial and final pressure ("\Delta V=V_0-V") and divide it by the initial value. Then multiply by 100% to express the answer in %:


"\epsilon=\frac{\Delta V}{V_0}\cdot100\%=\frac{V_0-V}{V_0}\cdot100\%,"

now it's time to substitute 0.926V0 for V:

"\epsilon=\frac{V_0-0.926V_0}{V_0}\cdot100\%=7.4\%."

Not 8%, but quite close.